OpenStudy (anonymous):

the line containing points (-1,3) and (3,8) has slope________ a) 22/7 b) 5/4 c)2.5 d)4/5

OpenStudy (itsbribro):

do you know the slope formula?

OpenStudy (itsbribro):

@Ammiel

OpenStudy (michele_laino):

please you have to apply this formula, in order to calculate the slope m: \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] where (x_1,y_1) and (x_2, y_2) are your points

OpenStudy (itsbribro):

well there is your formula lol

OpenStudy (michele_laino):

yes!

OpenStudy (itsbribro):

can you do the rest @Ammiel

OpenStudy (michele_laino):

please substitute the coordinates of your points @Ammiel

OpenStudy (michele_laino):

Hint: \[m=\frac{ 8-3 }{ 3-(-1) }=...\]

OpenStudy (anonymous):

hey @Michele_Laino and @itsbribro my internet failed sorry :/ i dont understand whats happening..

OpenStudy (anonymous):

is it 5div-3?

OpenStudy (anonymous):

*divided

OpenStudy (michele_laino):

plese, note thta: 3-(-1)=3+1=4, so?

OpenStudy (anonymous):

oh so then 5/4?

OpenStudy (michele_laino):

that's right!

OpenStudy (itsbribro):

good job :)

OpenStudy (anonymous):

:D yayyy thank you both of you :)

OpenStudy (itsbribro):

np :) i didnt help that much but yw

OpenStudy (michele_laino):

thanks! @itsbribro @Ammiel

OpenStudy (anonymous):

welcome :) and you did fine :)